\begin{thebibliography}{99}
\bibitem{Fujita:2011rd}
 T.~Fujita and N.~Kanda,
 ``A Proposal to Measure Photon-Photon Scattering,''
 arXiv:1106.0465 [hep-ph].

\bibitem{kanda}
N. Kanda, "Light-Light Scattering", to be published



\bibitem{euler}
  W.~Heisenberg and H.~Euler,
  ``Consequences of Dirac's theory of positrons,''
  Z.\ Phys.\  {\bf 98}, 714 (1936)

\bibitem{Moulin:1996vv}
 F.~Moulin, D.~Bernard and F.~Amiranoff,
 ``Photon-photon elastic scattering in the visible domain,''
 Z.\ Phys.\ C {\bf 72}, 607 (1996).

\bibitem{Bernard:2010dx}
 D.~Bernard {\it et al.},
``Search for Stimulated Photon-Photon Scattering in Vacuum,''
 Eur.\ Phys.\ J.\ D {\bf 10}, 141 (2000)
 [arXiv:1007.0104 [physics.optics]].

\bibitem{Moulin:2002ya}
 F.~Moulin and D.~Bernard,
 ``Four-wave interaction in gas and vacuum. Definition of a third order
 nonlinear effective susceptibility in vacuum,''
 Opt.\ Commun.\ {\bf 164} (1999) 137
 [arXiv:physics/0203069].

\bibitem{Lundstrom:2005za}
 E.~Lundstrom {\it et al.},
 ``Using high-power lasers for detection of elastic photon-photon scattering,''
 Phys.\ Rev.\ Lett.\ {\bf 96}, 083602 (2006)
 [arXiv:hep-ph/0510076].

\bibitem{Lundin:2006wu}
 J.~Lundin {\it et al.},
 ``Detection of elastic photon-photon scattering through four-wave mixing using high power lasers,''
 Phys.\ Rev.\ A {\bf 74}, 043821 (2006)
 [arXiv:hep-ph/0606136].

\bibitem{Kinoshita:2002ns}
 T.~Kinoshita and M.~Nio,
 ``Revised alpha**4 term of lepton g-2 from the Feynman diagrams containing an internal light-by-light scattering subdiagram,''
 Phys.\ Rev.\ Lett.\ {\bf 90}, 021803 (2003)
 [arXiv:hep-ph/0210322].

\bibitem{Jegerlehner:2009ry}
 F.~Jegerlehner and A.~Nyffeler,
 ``The Muon g-2,''
 Phys.\ Rept.\ {\bf 477}, 1 (2009)
 [arXiv:0902.3360 [hep-ph]].

\bibitem{Davier:2010nc}
 M.~Davier, A.~Hoecker, B.~Malaescu and Z.~Zhang,
 ``Reevaluation of the Hadronic Contributions to the Muon g-2 and to alpha(MZ),''
 Eur.\ Phys.\ J.\ C {\bf 71}, 1515 (2011)
 [arXiv:1010.4180 [hep-ph]].

\end{thebibliography}

\documentclass[showpacs]{revtex4}

\begin{document}

\preprint{arXiv:1106.0610v2 [hep-ph]}

\title{Comment on : A Proposal to Measure Photon-Photon Scattering}

\author{Denis Bernard}
\affiliation{Laboratoire Leprince-Ringuet, Ecole Polytechnique, CNRS/IN2P3, F-91128 Palaiseau, France }
 
\date{\today}%

\begin{abstract}
I comment on a recent preprint ``A Proposal to Measure Photon-Photon
Scattering'' that appeared recently as arXiv:1106.0465v1 [hep-ph].
\end{abstract}

\pacs{12.20.-m,42.55.-f}

\maketitle

The Authors of Ref. \cite{Fujita:2011rd} proposed that the low-energy
expression for the cross section of photon-elastic scattering
(Heisenberg and Euler \cite{euler}) : 

\begin{equation}
 {d\sigma\over d\Omega}\simeq {139\alpha^4\over (180\pi)^2 m^2} 
\left({\omega\over m} \right)^6 (3+\cos^2\theta )^2 
\end{equation}

, be replaced by (\cite{kanda}, to be published) : 

\begin{equation}
 {d\sigma\over d\Omega} \simeq {\alpha^4\over 
(12\pi)^2 \omega^2} (3+2\cos^2\theta +\cos^4\theta ) 
 \label{eq:kanda}
\end{equation}

where $\omega$ and $m$ denote the photon energy and the mass of
electron, respectively.
The orders of magnitude that were given  for $\omega \simeq 1$ eV are 
$2.3 \times 10^{-21}  \ \ {\rm cm}^2 $
and 
$9.3 \times 10^{-67} \ \ {\rm cm}^2 $, 
respectively.

~

Experimentally, an upper limit was  obtained first by colliding two
laser beams head-on, and searching for one of the scattered photons
(Moulin {\em et al.}  \cite{Moulin:1996vv}) :
\begin{equation}
\sigma <  10^{-39} \ \ {\rm cm}^2.  
\label{eq:direct}
\end{equation}

The limit was further improved with a stimulation of the reaction with
a third beam (Bernard {\em et al.} \cite{Bernard:2010dx}), based on
the computations of Moulin {\em et al.}  \cite{Moulin:2002ya} 
(that were revisited recently by Lundstrom, Lundin {\em et al.}
\cite{Lundstrom:2005za,Lundin:2006wu}).

\begin{equation}
\sigma < 1.5 \times 10^{-48} \ \ {\rm cm}^2. 
\label{eq:stimulated}
\end{equation}

These two experimental publications
\cite{Moulin:1996vv,Bernard:2010dx} are indeed mentionned in
Ref. \cite{Fujita:2011rd}, but it would be interesting to understand
how the prediction of eq. (\ref{eq:kanda}) can face the upper limits
of eqs. (\ref{eq:direct}) and (\ref{eq:stimulated}).

~

Also, it should be noted that the QED computation of
the 4-photon loop diagrams  (See, e.g.,
\cite{Kinoshita:2002ns} and a review at \cite{Jegerlehner:2009ry}) take part into the prediction of the lepton
``anomalous''  magnetic moment, which is found to be in fair (i.e., within three
standard deviations) agreement with the experimental measurement (e.g.
\cite{Davier:2010nc}).

\end{document}